# Deep learning architecture for decrypting information on the event horizon

### v7

@@ Abstract
According to βt Hooft, to recover the invariance under the PoincarΓ© group in a holographic setting, the evolution law for the direction orthogonal to the given surface and the time evolution law must commute. The condition of commutativity assumes that the time-evolution law on the given surface and the target surface is the same. Meanwhile, the AdS/CFT correspondence implies that there exists a mapping from the black hole horizon to the conformal boundary. In this paper, it is conjectured that in a large AdS black hole in thermal equilibrium, this mapping and the time-evolution law will commute using the properties derived from the infinite volume limit. Also, we provide a computational model in the context of AdS/CFT correspondance assuming this commutativity in a holographic setting. Specifically, we will use deep learning algorithms to find a model that tells us how pieces of information on the event horizon of an AdS black hole are decrypted into conformal boundary. By combining commutativity with AdS/CFT, we might move beyond the view that the conformal boundary implements the holographic screen and find clues about how information pieces on the event horizon are transformed into information pieces on the conformal boundary.
1 Introduction
@@
The holographic principle is considered a fundamental characteristic of quantum gravity theory. This principle started from the fact that the entropy of a black hole is proportional to the area of the horizon and was extended to the general case[1]. Considering unitarity, entropy and counting arguments, it was pointed out that the 2-dimensional surface should determine all information in the 3+1 dimension. Note that considering the limit where the two-dimensional space becomes flat, this means that there is a transform in the direction perpendicular to that flat space. Furthermore, since the ultimate theory will satisfy PoincarΓ© invariance, this transformation and the time evolution will commute with each other. The proposed cellular automaton model, in which information in a two-dimensional space determines all information in the 3+1 dimension, had only a linear structure and could not include interactions. Therefore, we aim to build a non-linear model from an AdS/CFT perspective.
@@
The flat space-time used to construct this model was implemented as a holographic screen, and its connection with string theory was revealed[2]. For the event horizon, it can actually be mapped by following light rays. Here, the focusing theorem guarantees that there exists at least one map satisfying the holographic principle. This mapping unfolds the entropy on the black holeβs horizon onto flat surface, so black hole physics still lives on this screen. However, since it does not evolve according to Lorentzian time on a holographic screen, it is difficult to define a cellular automaton model on this screen[3]. Moreover, since the degree of freedom is not generally preserved, it is difficult to expect it to follow conventional physics. Instead, we will assume that the local nature of physics allows us to select small pieces of information from sufficiently large horizon and conformal boundary respectively, and based on this assumption, an approximation method that flattens the selected information pieces will be adopted.
@@
In the AdS/CFT correspondence, the conformal boundary is realized as a holographic screen[4]. This correspondence guarantees the existence of a transformation that connects the event horizon and the conformal boundary. The assumption of commutativity is a very strong one, thus finding a model that satisfies this property in a typical AdS/CFT setting is inevitably very limited. Therefore, we will only consider the thermal equilibrium state of the large AdS black hole. In the context of the black hole paradox, the piece of information inscribed on the horizon can be considered to be the encrypted information of the conformal boundary[5]. From this perspective, the function that goes from the horizon to the conformal boundary will be called the decrypting function. Here, saying that this function is reversible is equivalent to saying that no information is lost in a black hole. In this paper, we propose a computational model that can find this decrypting function with minimal physical constraints.
@@
These results imply that deep learning can be an effective computational tool even in quantum gravity. Therefore, the proposed computational model adopts a deep learnig algorithm.
2 Commutativity and AdS/CFT correspondonce
The object to be dealt with here is the AdS black hole, so the many-body problem will be considered. Additionally, for the purpose of building the computational model of AdS/CFT, we will assume that a time evolution law of microstates forming an ensemble exists as in the cellular automaton(CA) interpretation, but we will exclude fundamental discussions of quantum mechanics such as the Copenhagen interpretation[14]. Furthermore, we will assume that locality is satisfied.
The holographic principle implies the existence of an evolution law in the holographic direction. This law will be called the operator. Assuming that the invariance under the PoincarΓ© group is recovered in the ultimate holographic theory, the time evolution law and must commute[1]. Furthermore, to assume this invariance, the spaces considered must be flat. Let us first consider an approximation for flatness in typical AdS/CFT setting with block hole. In this setting, the conformal boundary is considered as the spacetime where the information on horizon is decrypted. To meet requirement for flatness, we will use the fact that the conformal boundary becomes locally flat when the black holeβs mass becomes infinitely large[15]. In other words, it has the same effect as taking an inifinite volume limit. Then the boundary becomes that is flat. Note that the physics above this conformal boundary is described by gauge theory, so it is different from the holographic screen. Also, for the event horizon, as the mass of the black hole increases, the radius of the horizon can be infinitely increased. Therefore, we can make the conformal boundary and horizon as large as we want. If we select sufficiently small information pieces from a very large surface, we can say that these pieces are approximately flat.
Additionally, for large AdS black hole, the radius of the horizon is much larger than the radius of curvature of the AdS in this limit. Therefore, if the size of the black hole becomes very large but is finite, the entropy above the conformal boundary and horizon is proportional. Moreover, we can expect the entropies to be equal at an appropriate βt Hooft parameter . Therefore, to compare the same amount of entropy, finite scale will be considered in the proposed model.
Combining the two preceding conditions, the proposed model considers a sufficiently large but finite volume of the boundary. Indeed, due to locality, it can be expected that such information pieces exist on each surface and that they will be simply connected. These selected pieces of information are related by transform. With this setting, a more general form of commutativity can be written as Equation 1, where the our assumption says that and .
Here, and are the time evolution law at the conformal boundary of AdS space and event horizon of large AdS black hole, respectively. They can be simply denoted as if .
To find a specific model that satisfies and in 1, consider the large Schwarzschild black hole solution in space. In this case, it is known that gauge theory above the conformal boundary has high temperature behavior[15]. Also, the corresponding AdS black hole horizon can reach thermal equilibrium[16]. Therefore, any piece of information on the conformal boundary and the corresponding piece of information on the event horizon can be considered as one possible microstate. Here, the change over time in the piece of information inscribed on the horizon is due to quantum fluctuations near the horizon. The Hawking temperature formula tells us the rate at which quantum fluctuations occur, and since this is expressed in coordinate time, pieces of information also change according to coordinate time. Additionally, after the black hole in AdS reaches thermal equilibrium, the emission and absorption rate must be the same, so the absorption rate also changes depending on coordinate time. Specifically, the rate at which Hawking-radiated particles are reflected at the boundary and re-absorbed has the same magnitude and opposite sign as the emissivity due to Hawking radiation[17]. Therefore, coordinate time can be given as a time parameter for the microscopic state on the event horizon same as on the conformal boundary. Also, the AdS/CFT correspondence predicts that a state corresponding to the quantum fluctuations on the horizon also exists on the conformal boundary. Furthermore, since it is expected that PoincarΓ© invariance will be restored at the Planck scale, it is conjectured that the operator and the operator for information pieces on these boundaries will commute. Note that the duality can be confirmed more clearly through deformation by adding branes to given space[18]. This might be considered a deformation where is maintained but becomes a value other than 1 in Equation 1.
@@ 2 Commutativity and AdS/CFT correspondonce
Now, considering all these points in this section, as shown in Figure1 a), commutativity means that the results of time evolution at the event horizon and then decrypting to a conformal boundary are the same as those of time evolution after decrypting to a conformal boundary.
3 Deep learning for holographic system
Assuming that we know the time evolution as in CA, let us implement a computational model to find the operator using a deep learning algorithm. In other words, assuming that is given, will be trained as in Figure1 b). The spatial dimension of the event horizon corresponds to the spatial dimension of the conformal boundary. For visualization purposes, we consider here an AdS/CFT setting with one spatial dimension reduced, i.e., where the spatial dimensions of the boundary of AdS space and the horizon of AdS black hole are both 2. It is straightforward to recover that dimension in proposed models. Additionally, for simplicity, classical CA will be considered, but the full description must be quantum mechanical, and therefore the computational model should also adopt a quantum mechanical deep learning algorithm.
@@ 3 Deep learning for holographic system
To implement the proposed computational model, it is crucial to note that in deep learning algorithms, updating weights occur with each batch. Consequently, labels will be generated using the intermediate state of the model updated up to the given batch point. Therefore, as illustrated in the proposed algorithm below1, the loops should be decomposed into those for epochs and those for batches.
@@
In the above proposed algorithm, various functions are employed to facilitate the process. IntermediateModel is a model trained up to the current batch point. By placing intermediate states in a loop, learning is possible recursively. GenerateRandomBinaryTensor function generates raw random data to be used as input for the model. PreprocessData formats raw data to make it suitable for model input, ensuring it is appropriately normalized and transformed. GetBatch function extracts a specific batch from the entire dataset, defined by the batch size, and uses the given batch number and size to partition the data, returning the specified batch. LogitToPrediction converts the modelβs output, known as logits, into final predictions. This conversion is typically achieved through softmax or sigmoid functions, which transform logits into probability values that serve as the modelβs predictions. ConvertToOneHot transforms the given label data into one-hot encoded vectors, which are essential for multi-class classification tasks. Finally, TrainModel uses the provided input and target data to train the model, adjusting the model parameters to minimize the loss function throughout the training process.
@@
For the operator assumed above, at least the following two properties must be satisfied. The first condition is the locality assumed earlier. This is satisfied if the current cell is determined by the neighboring cells in the specified bound like Conwayβs game of life. Second, roughly speaking, to preserve quantum mechanical probability it should be time-reversible. That is, the cells in the past should be able to be determined from current cells. Since two consecutive CA layers are needed to define the rule going to the future or the past, the evolution law for one time step can be split into a two-step procedure. Here, the block CA [19] will be adopted as an example of time-evolution law. The well-known block CA rules are as follows.
@@
For edge cells after the first rule is applied, it is assumed that cells on opposite sides are grouped together to form a torus. It is known that it satisfies locality by definition and is time-reversible. Note that when we know the rules for even and odd sites like block CA, it is known that Hamiltonian can be expressed as an infinite sequence of commutators of these rules using BakerβCampbellβHausdorff formula[14].
On the other hand, we know that the block CA rule must satisfy linearity to become the time-evolution law of the quantum system. The fact that the linearity is satisfied can be seen easily by showing that the block CA rule is expressed as a matrix multiplication. This matrix is block diagonal and the elements of this block diagonal matrix are functions that, given initial conditions, become 0 or 1 over time. If the block CA were a real quantum system, each element would be a time-dependent function determined by SchrΓΆdinger equation, Hamiltonian, and initial condition.
@@
As a dataset, random input will be used in this paper. A random dataset is defined on a event horizon, where the values 0βs and 1βs on each lattice might indicate the absence or presence of a photon at that position, respectively. After mapping information to the lattice like this example, the operator converts it to information in the corresponding conformal boundary. Although discussions on how to encode physical states into information using bits as the unit are crucial in the proposed computational model, it is not the objective of this paper.
Now, the architecture of CNN to learn operater is presented in Figure2.
@@
Here, a convolution-deconvolution[20] architecture was used recursively. The proposed network follows a conventional deep learning architecture used in image segmentation, characterized by a convolution-max pooling-deconvolution block. The depth of these blocks is considered irrelevant to the current problem, as the correlation between the depth and the loss is not discernible critically. However, dropout layers and regularization are critically important for the problem we aim to solve. This is because, among the solutions that satisfy commutativity, there are cases where is a constant function or trivial solutions. Without these techniques, the model tends to learn trivial solutions.
@@
However, more importantly, it is crucial to carefully select the loss function. As mentioned earlier, the proposed model tends to learn trivial solutions, so to avoid this, the loss function must be customized to penalize trivial solutions. In this model, we added a conditional statement to assign a very large loss value when a trivial solution is encountered, thereby encouraging the learning of non-trivial solutions. Nonetheless, non-trivial solutions are not necessarily physically meaningful. To obtain physically meaningful solutions, the model must incorporate the requirements derived from the holographic principle in the context of AdS/CFT. For example, among the numerous solutions that can take, the selected solution often depends on the weight initialization. For more realistic learning, it is essential to initialize the weight based on a physically correct interpretation of the weight[11].

### v8

# Commutative Evolution Laws in Holographic Cellular Automata: AdS/CFT, Near-Extremal D3-Branes, and a Deep Learning Approach

@@ Abstract
According to ’t Hooft, restoring Poincaré invariance in a holographic cellular automaton (CA) requires two distinct evolution laws that commute. We explore how this is realized in the AdS/CFT framework, assuming commutativity as a fundamental principle—much like general covariance once did—for encoding curvature. In our setup, physical processes in a given spacetime are encoded in a CA; to preserve Poincaré symmetry, the spacetime curvature must effectively vanish, so we consider a near-extremal black D3-brane solution, in which both the stretched horizon and the conformal boundary are approximated by Minkowski space. AdS/CFT implies a spatial evolution law connecting these hypersurfaces. Commutativity means the final state does not depend on the order of time evolution on each hypersurface and spatial evolution between them, forcing the time evolution law on the horizon and boundary to coincide. To satisfy all these conditions, we aim to demonstrate that the spatial evolution law inevitably encapsulates the curvature of the bulk, including quantum effects. For a computational model, we compactify the hyperplanes to tori, reducing the degrees of freedom to a finite number; taking these tori to infinite size then restores Poincaré symmetry. We propose a deep learning algorithm that, given a known time evolution law and commutativity, deduces the corresponding spatial evolution law.
I Introduction
@@
The holographic principle is regarded as a fundamental aspect of quantum gravity theory. This principle started from the fact that the entropy of a black hole is proportional to the area of the horizon and was extended to the general case[1]. Considering unitarity, entropy and counting arguments, it was pointed out that the 2-dimensional surface should determine all information in the 3+1 dimension. Note that considering the limit where the two-dimensional space becomes flat, this means that there is a transform in the direction perpendicular to that flat space. Furthermore, since the ultimate theory will satisfy Poincaré invariance, this transformation and the time evolution will commute with each other. The proposed cellular automaton(CA) model, in which information in a two-dimensional space determines all information in the 3+1 dimension, had only a linear structure and could not include interactions. Therefore, we aim to build a non-linear model from an AdS/CFT perspective.
@@
The flat space-time used to construct this model was implemented as a holographic screen, and its connection with string theory was revealed[2]. An event horizon can in fact be mapped by following light rays. Here, the focusing theorem guarantees that there exists at least one map satisfying the holographic principle. This mapping unfolds the entropy on the black hole’s horizon onto a flat surface, allowing black hole physics to remain on this screen. However, because it does not evolve under Lorentzian time on a holographic screen, defining a CA model on this screen becomes difficult[3]. Moreover, since the degree of freedom is not generally preserved, it is difficult to expect it to follow conventional physics. Finally, in the AdS/CFT correspondence, the conformal boundary is realized as a holographic screen[4][5]. Also, this correspondence guarantees the existence of a transformation that connects the event horizon and the conformal boundary.
We aim to take a step further by combining this remarkable fact with the commutative property and the CA model. Specifically, considering the AdS/CFT correspondence, the main objective will be to explore how the argument used in ’t Hooft’s original paper[1] can be justified and to interpret the significance of the commutative property. The motivation for focusing on commutativity in this paper is derived from ’t Hooft’s CA paradigm and is not identical to the standard renormalization group (RG) interpretation. When two operations commute, their final outcome is independent of the order in which they are applied, thereby preventing path-dependent loss or distortion of information. In quantum mechanics, the uniqueness of mapping one initial state to a single final state precisely characterizes unitarity (i.e., information conservation). Hence, a conceptual link emerges such that “commutativity implies information preservation (unitarity).” Therefore, positing commutativity can be regarded as a plausible assumption necessary for maintaining information conservation. However, this should not be taken to equate strictly to quantum-mechanical unitarity. True unitarity entails a norm-preserving linear operator on a Hilbert space. Our commutativity assumption is instead a toy-model principle for preventing path-dependent loss of information, rather than a complete statement of quantum unitarity. Since defining the commutation relation on the boundary of a given spacetime using a CA model requires the boundary surfaces to be flat, we consider a near-extremal black D3-brane solution. Accordingly, we will propose a possible interpretation about the commutation relations in AdS space, considering the thermal equilibrium state of a near-extremal black D3-brane. One of the virtues of this holographic model is that it allows us to disregard the curvature within the given bulk. Paradoxically, we will find that the transformation defined in the holographic direction within this interpretation rather encodes information about that curvature.
@@
In the context of the black hole paradox, the piece of information inscribed on the horizon can be considered to be the encrypted information of the conformal boundary[7]. From this perspective, the function that goes from the horizon to the conformal boundary will be called the decrypting function. Here, saying that this function is reversible is equivalent to saying that no information is lost in a black hole. In this paper, we propose a computational model that can find this decrypting function.
@@
These results imply that deep learning can be an effective computational tool even in quantum gravity. Therefore, the proposed computational model adopts a deep learning algorithm.
II Commutativity and AdS/CFT correspondence
Here, we focus on the AdS black brane, thus bringing a many-body problem into consideration. In this case, it is known that gauge theory above the conformal boundary has high temperature behavior[4][6][19]. Also, the corresponding black brane horizon can reach thermal equilibrium. Therefore, any piece of information on a given boundary can be regarded as a microstate that evolves according to the time-evolution law while preserving degrees of freedom. Additionally, for the construction of our computational AdS/CFT model, we assume that the microstates forming an ensemble are encoded in a cellular automaton, thus following the logic of the cellular automaton interpretation, but we will exclude fundamental discussions of quantum mechanics such as the Copenhagen interpretation or any other interpretational aspects of quantum mechanics[18]. The holographic principle implies the existence of an evolution law in the holographic direction. This law will be called the operator. Assuming that the invariance under the Poincaré group is recovered in the ultimate holographic theory, a certain time evolution law and must commute[1]. We will compare the information defined on the spatial slices of two spacetimes connected by .
To encode information about physical processes in a cellular automaton (CA) and maintain Poincaré symmetry, the spacetime under consideration must be Minkowski. Otherwise, it becomes difficult to define a time evolution law by slicing the spacetime into equal-time surfaces. Moreover, Poincaré symmetry itself is originally defined in Minkowski space. In the context of AdS/CFT, two Minkowski hypersurfaces can be found in the near-extremal D3-brane solution in decoupling limit. Maldacena’s near-extremal D3-brane metric (often referred to as Eq. (2.5) in [4]) is:
where is the string coupling, is the number of D3-branes, and sets the horizon scale. We assume so that the supergravity approximation remains valid. To see how the metric becomes approximately Minkowski near a stretched horizon, consider fixing the radial coordinate at
Since there is no radial variation on this slice, the term involving does not contribute to the induced metric. Separately, we expand
and substitute into (2.5). At leading order in , the part simplifies to
revealing a Minkowski-like form in the directions, up to an overall scale factor. Meanwhile, in the limit , we have
and the metric becomes
which becomes conformally equivalent to flat Minkowski space in the directions. It is worth emphasizing that the approximate Minkowski form at the stretched horizon and at the asymptotic boundary applies only within these respective local patches. The bulk itself still possesses nontrivial curvature, and these Minkowski regions do not imply flatness throughout the entire spacetime. The time delay (or gravitational redshift) we will discuss arises because information must traverse a curved bulk region, despite the local near-flatness at the two boundaries.
Remark on Microstate Dynamics. Even though is treated as a constant for this near-horizon analysis (i.e., a stationary black brane), the underlying microstates associated with the brane can still evolve in time. In a holographic or cellular automaton (CA) interpretation, one may encode these microstates into configurations that change while the macro-scale parameter remains fixed overall, maintaining thermal equilibrium. Hence, the bulk geometry at the stretched horizon can be nearly Minkowskian, yet the microscopic degrees of freedom can undergo nontrivial dynamics (e.g., Hawking-like processes balanced by an external bath), preserving the equilibrium at the macroscopic level.
On the other hand, when considering these two spacetimes together, we define the time variable merely as an index for a collection of space-like surfaces. Since the curvature of spacetime does not depend on the choice of coordinates, both spacetimes remain flat. To delve deeper, consider a particular physical process at the stretched horizon. Information on the stretched horizon corresponding to that process can be transferred to the conformal boundary through Hawking radiation. Maybe, if certain particle falls into the horizon, this information can also be transferred by particles produced through quantum fluctuations, reaching the conformal boundary. Regardless of the mechanism, an observer at the conformal boundary will experience time dilation due to gravity. However, for commutativity to hold in our assumed model, the law of time evolution must remain identical in both spacetime frames. Let us denote this time evolution law by . Therefore, to satisfy these conditions, the time delay after a spatial translation should not be expressed in the form of , but rather, it should already be incorporated into . Although it may be hard to imagine, the CA model can produce a vast number of possible evolutions based on each piece of information, so it is reasonable to expect that there exist pairs of and satisfying these conditions.
Let us return to the issue of information on the stretched horizon being transmitted to the conformal boundary. Note that a length scale is not defined on the conformal boundary. By first matching the areas of the stretched horizon and the conformal boundary, we establish a common length scale for both planes. Next, consider a pair of short-wavelength photons produced in the bulk; one travels to the horizon, and the other to the boundary. Because each photon’s wavelength is sufficiently short, it can arrive at specific locations while carrying multiple bits of information—thus making the impact of the bulk’s curvature more transparent. In other words, the curved paths they follow lead them to distinct, asymmetrical arrival points, vividly illustrating how the bulk is distorted. In this scheme, the evolution law —which transfers information from the CA at the stretched horizon to the boundary—encodes the resulting displacement as a permutation of states. As a result, the difference in each photon’s endpoint, driven by the bulk curvature, is captured within in the form of a rearrangement, effectively mapping the curvature into a permutation structure.
However, as highlighted in discussions of entanglement entropy, the AdS/CFT correspondence inherently features nonlocality[9]. This nonlocality, which arises from quantum mechanical effects, is also expected to permute or spread information beyond the classical distortions of time and space mentioned earlier. In fact, without this nonlocality, any CA-based discussion would likely have been rendered meaningless. For example, consider a metric that exhibits a time delay but no spatial distortion. In the absence of quantum effects, a CA encoding the information of the stretched horizon would be mapped identically onto the conformal boundary. However, assuming there is a time delay, the time evolution on the conformal boundary should remain consistent with that of the stretched horizon while also appearing to reflect this delay. Yet, because the CAs themselves are entirely identical, its evolution would also remain the same, ultimately making the time delay disappear. Therefore, nonlocality is an essential property that must be addressed alongside the CA model.
@@ II Commutativity and AdS/CFT correspondence
Now, in the context of AdS/CFT, a possible interpretation of the commutation relation is that the operator and the operator on the stretched horizon commute with each other. Considering all the points so far, as shown in Figure1 a), it means that the results of time evolution at the stretched horizon and then decrypting to a conformal boundary are the same as those of time evolution after decrypting to a conformal boundary.
Moreover, the previously discussed requirement that the time delay and spatial distortion be reflected in the function implies that the function contains information about the curvature of the AdS spacetime. This fact is represented in Figure1 a) also, where the bulk interval is shown to align with the location of the function . This suggests that, just as general covariance served as a guide in formulating equations to compute curvature in classical theories, commutation relations could serve as a guide in constructing models for computing curvature with quantum effects in a theory of quantum gravity.
III A Deep Learning Algorithm for a Holographic System
Let us now consider the problem of constructing a computational model that satisfies commutativity. In order to build such a model, as in the original work[1], the amount of information defined on each hyperplane must remain finite. A straightforward and intuitive way to maintain finite entropy, while preserving zero curvature in Minkowski space, is to compactify the space into a torus. Implementing this toroidal compactification on the black brane metric (cf. Eq. (2.5)) means imposing periodic boundary conditions, expressed via the following identifications:
(etc. for any additional coordinates). |
thus effectively turning the black brane into a torus. However, this compactification breaks the continuous Poincaré symmetry down to a discrete one. Note that Poincaré symmetry is not a necessary and sufficient condition for commutativity. As mentioned earlier, commutativity is an assumed condition from the perspective of information conservation, whereas Poincaré symmetry has been treated merely as a necessary condition for constructing the model in a straightforward manner. If commutativity is truly a universal property rather than a feature confined to a particular physical system, then it must continue to hold even if Poincaré symmetry is lost, i.e. even when the spacetime is compactified to a torus. Moreover, discussions of AdS/CFT and thermodynamics still remain valid, supported by the general arguments for topological AdS black holes[8]. Although the zero curvature still allows us to apply a CA model, the boundary conditions in the and directions (for instance) introduce slight modifications to the paired evolution laws. Nonetheless, if we take the limit where the torus sizes go to infinity, Poincaré symmetry will be restored, enabling us to continue the previous line of reasoning without significant difficulty.
On the other hand, one might consider omitting compactification and instead restricting attention to a finite region of an infinitely large Minkowski spacetime, thereby ostensibly maintaining the symmetry locally while keeping the degrees of freedom finite. For instance, in the Ryu–Takayanagi prescription for the entanglement entropy of an AdS black hole[9], one can choose a minimal surface that wraps the horizon partially, thereby relating the thermodynamic entropy at the horizon to the corresponding entanglement entropy in the CFT. Yet, because the entanglement entropy in a CFT is ultimately defined by the entanglement between a chosen region and its complement, it becomes challenging to quantify how local information at the horizon might be nonlocally distributed across the conformal boundary. Consequently, one must consider the entire conformal boundary, at which point global curvature arises and the situation becomes significantly more complicated.
Therefore, considering a setup that includes compactification, and assuming we know the time evolution as in a CA, let us implement a computational model to find the operator using a deep learning algorithm. In other words, assuming that is given, will be trained as in Figure1 b). The spatial dimension of the event horizon corresponds to the spatial dimension of the conformal boundary. For visualization purposes, we represent the physical processes on two 3-dimensional spatial hypersurfaces encoded in the CA as 2-dimensional grids. Additionally, due to the correspondence, the entropy is the same, so we consider 2-dimensional grids of identical size. For simplicity, classical CA will be considered. Although we employ a classical (block) CA as a toy model, one should note the fundamental differences between Boolean-state cellular automata and genuine quantum systems. In a quantum theory, states reside in a complex Hilbert space and evolve unitarily, thereby preserving norms and accommodating entanglement. By contrast, classical CA models typically involve discrete 0/1 updates without capturing complex amplitudes. While this paper draws upon CA-based intuition, fully realizing quantum features would require a more sophisticated, genuinely quantum cellular automaton framework. However, note that nonlocality stemming from quantum effects manifests as the mixing and spreading of information, and it can be partially emulated even within classical CA models. Note that our discussion of nonlocality concerns how quantum effects can spread or mix information in ways that exceed classical correlations. Meanwhile, in a classical CA, ‘nonlocal’ rearrangements often amount to permutations of discrete states. Although useful for illustrating certain aspects of holography, these permutations cannot fully replicate the physics of quantum entanglement. Thus, we caution readers that our classical CA remains an approximation, and genuine quantum nonlocality requires a more rigorous operator-based or entanglement-based formulation.
@@ III A Deep Learning Algorithm for a Holographic System
The computational model proposed in Figure1 b) corresponds to supervised learning in the domain of deep learning, but it can be observed that it exhibits a recursive structure where intermediate results during training serve as labels that act as ground truth in supervised learning. To implement the proposed computational model, it is crucial to note that in deep learning algorithms, updating weights occur with each batch. Consequently, labels will be generated using the intermediate state of the model updated up to the given batch point. Therefore, as illustrated in the proposed algorithm below Figure2, the loops should be divided into epoch loops and batch loops.
@@
In the proposed algorithm, various functions are employed to streamline the training process. IntermediateModel refers to a model that has been trained up to the current batch, thereby allowing recursive learning when these intermediate states are placed inside a loop. The GenerateRandomBinaryTensor function produces raw random data to be used as model input. Meanwhile, PreprocessData ensures that this raw data is suitably normalized and transformed for the model. The GetBatch function extracts a specific batch from the dataset — defined by the batch number and size — by partitioning the data accordingly. The LogitToPrediction function converts the model’s outputs (logits) into final predictions. This is typically achieved via softmax or sigmoid functions, which transform logits into probability values. ConvertToOneHot then encodes the label data into one-hot vectors, a common practice for multi-class classification. Finally, TrainModel takes both the input and target data to update the model’s parameters, minimizing the loss function over the course of training.
@@
For the operator assumed above, at least the following two properties must be satisfied. The first condition is locality, as in a typical CA model (to be distinguished from the nonlocality in AdS/CFT). This is satisfied if the current cell is determined by the neighboring cells in the specified bound like Conway’s game of life. Second, roughly speaking, to preserve quantum mechanical probability it should be time-reversible. That is, the cells in the past should be able to be determined from current cells. Since two consecutive CA layers are needed to define the rule going to the future or the past, the evolution law for one time step can be split into a two-step procedure. Here, the block CA [20] will be adopted as an example of time-evolution law. The well-known block CA rules are as follows.
@@
For edge cells after the first rule is applied, it is assumed that cells on opposite sides are grouped together to form a torus to align with the compactification condition. It is known that it satisfies locality by definition and is time-reversible. Note that when we know the rules for even and odd sites like block CA, it is known that Hamiltonian can be expressed as an infinite sequence of commutators of these rules using Baker–Campbell–Hausdorff formula[18].
On the other hand, we note that in order for the block CA rule to serve as a toy model of quantum time evolution, it should exhibit (at least approximate) linearity. We can verify this linearity by showing that the block CA rule can be written in terms of a matrix multiplication. In this toy model, the matrix is block diagonal, and although its elements can take binary values (0 or 1) at each discrete time step, we emphasize that each element is effectively time-dependent through iterative updates rather than remaining a fixed constant. Of course, in a genuine quantum system, each matrix element would follow a continuous, complex-valued time evolution as governed by the Schrödinger equation, the Hamiltonian, and the initial condition, rather than collapsing to 0 or 1. In this sense, the “linearity” of a classical CA is restricted to a local rule over , which remains fundamentally distinct from a unitary operator on a complex Hilbert space in a quantum system.
@@
As a dataset, random input will be used in this paper. A random dataset is defined on a stretched horizon, where the values 0’s and 1’s on each lattice might indicate the absence or presence of a photon at that position, respectively. After mapping information to the lattice like this example, the operator converts it to information in the corresponding conformal boundary. Although discussions on how to encode physical states into information using bits as the unit are crucial in the proposed computational model, it is not the objective of this paper.
We now present the CNN architecture used to learn the operator in Figure4.
@@
Here, a convolution-deconvolution[21] architecture was used recursively. The proposed network follows a conventional deep learning architecture used in image segmentation, characterized by a convolution-max pooling-deconvolution block. The depth of these blocks is considered irrelevant to the current problem, as the correlation between the depth and the loss is not discernible critically. However, dropout layers and regularization are critically important for the problem we aim to solve. This is because, among the solutions that satisfy commutativity, there are cases where is a constant function or trivial solutions. Physically, these trivial solutions do not represent meaningful transformations in the AdS/CFT context. During the experimental process, it was observed that the model tended to learn trivial solutions in the absence of these techniques.
@@
However, more importantly, it is crucial to carefully select the loss function. As mentioned earlier, the proposed model tends to learn trivial solutions, so to avoid this, the loss function must be customized to penalize trivial solutions. In this model, we added a conditional statement to assign a very large loss value when a trivial solution is encountered, thereby encouraging the learning of non-trivial solutions. Nonetheless, non-trivial solutions are not necessarily physically meaningful. To obtain physically meaningful solutions, the model must incorporate the requirements derived from the holographic principle in the context of AdS/CFT. For example, among the numerous solutions that can take, the selected solution often depends on the weight initialization. For more realistic learning, it is essential to initialize the weight based on a physically correct interpretation of the weight[15]. For instance, according to the interpretation of presented earlier, the initialized weights could correspond to the classical solution of general relativity. Then, the training can be interpreted as a correction due to quantum effects corresponding to nonlocality. In a future, fully quantum framework, this approach could be extended by embedding unitarity or entanglement constraints directly into the architecture (e.g., through specialized layers or loss functions).